OpenStudy (anonymous):

Solve the following equation: 3x+1/3x+2 = x+2/x-5 A. -9/22 B. -9/23 C. -3/8 Please and thanks!!

hero (hero):

\[\frac{3x + 1}{3x + 2} = \frac{x + 2}{x - 5}\] \[\frac{3x + 2 - 1}{3x + 2} = \frac{x - 5 + 7}{x - 5}\] \[\frac{3x + 2}{3x + 2} - \frac{1}{3x + 2}= \frac{x - 5}{x - 5} + \frac{7}{x - 5}\] \[1 - \frac{1}{3x + 2}= 1 + \frac{7}{x - 5}\] \[ - \frac{1}{3x + 2}= \frac{7}{x - 5}\] \[-1(x - 5) = 7(3x + 2)\] \[5 - x = 21x + 14\] \[-(14 - 5) = (21 + 1)x\] \[-9 = 22x\] \[-\frac{9}{22} = x\]